OpenStudy (anonymous):

if the 1st term and the common diff of a G.P are each equal to x + sqrt(5-x^2) i then the modulus of its nth term is?

OpenStudy (anonymous):

\[|x + \sqrt{5-x^2} i |=\sqrt{x^2+5-x^2}=\sqrt{5}\]

OpenStudy (anonymous):

so make a G.P of modulus first term sqrt{5} common diff sqrt{5}

OpenStudy (anonymous):

thxx